# Moral Susceptibility and Robustness under Persona Role-Play in Large Language Models

### v1

@@ 1 Introduction
As large language models (LLMs) move into interactive, multi-agent settings, reliable benchmarks for their social reasoning are essential. Recent evaluations probe theory-of-mind, multi-agent interactions under asymmetric information, cooperation, and deception through controlled role-play and game-theoretic tasks [zhou2024sotopia, pan2023machiavelli, bianchi2024negotiationarena, chen-etal-2024-tombench, costa2025deceivedetectdiscloselarge]. Complementary datasets benchmark social commonsense, moral judgment, and self-recognition capabilities [sap-etal-2019-social, hendrycks2021ethics, bai2025knowthyselfincapabilityimplications]. Motivated by this landscape, we focus on moral judgment as a core facet of social decision-making and alignment.
This paper introduces a benchmark that combines persona role-play—prompting a LLM to assume a specific character—with the Moral Foundations Questionnaire [moralfoundations2017questionnaires], a widely used instrument in moral psychology that measures five moral foundations: Harm/Care, Fairness/Reciprocity, In-group/Loyalty, Authority/Respect, and Purity/Sanctity [graham2009liberals, haidt2007when, moralfoundations2017questionnaires]. We elicit LLMs to respond to the MFQ while role-playing personas drawn from ge2025scalingsyntheticdatacreation. From these responses, we define two complementary quantities: moral robustness, the stability of MFQ scores over personas under repeated sampling, and moral susceptibility, the sensitivity of MFQ scores to persona variation. See Fig. 1 for a conceptual overview diagram. These metrics are defined in Eq. (4) and Eq. (7), each with foundation-level decompositions and uncertainty estimates.
@@
Applying this framework across contemporary model families and sizes, we find that model family accounts for most of the variance in moral robustness, with no systematic effect of model size. In contrast, moral susceptibility shows a mild family effect but a clear within-family size trend, with larger variants being more susceptible. Among individual models, Claude 4.5 Sonnet is the most robust and Grok 4 Fast the least. Conversely, Gemini 2.5 Flash is the most susceptible, while GPT-5 Nano is the least. Overall, we observe a a positive correlation between robustness and susceptibility, which is more pronounced at the family level, as seen in Section 3.3.
Recent research has examined the moral and social behavior of LLMs through the lens of the MFQ, exploring their value orientations, cultural variability, and alignment with human moral judgments [abdulhai-etal-2024-moral, nunes2024hypocrites, aksoy2024whose, bajpai2024insights, ji2025moralbenchmoralevaluationllms]. Parallel efforts study persona role-playing as a mechanism for conditioning model behavior, including benchmarks, interactive environments, and diagnostic analyses [tseng2024two, wang2023rolellm, samuel2025personagym, yu2025rpgbench, xu2024character, elboudouri2025rpeval, bai2025concept]. Our MFQ persona framework bridges these directions by systematically quantifying how persona conditioning alters moral judgments, separating the effects of repeated sampling (moral robustness) from those of persona variation (moral susceptibility). In addition, we report MFQ profiles for both unconditioned and persona-conditioned settings, providing a comparative view of baseline moral tendencies and persona-driven moral shifts across models.
@@ 2.1 Moral Foundation Questionnaire
The Moral Foundations Questionnaire [moralfoundations2017questionnaires] is a widely used instrument in moral psychology [graham2009liberals, haidt2007when, moralfoundations2017questionnaires] and comprises 30 questions split into two sections. The first includes 15 relevance judgments, which assess how relevant certain considerations are when deciding what is right or wrong, and the second includes 15 agreement statements, which measure the level of agreement with specific moral propositions [graham2011mfq, moralfoundations2017questionnaires]. In both sections, respondents answer each item using an integer scale from 0 to 5, representing in the first section the perceived relevance of the consideration and in the second the degree of agreement with the statement (see Appendix A for a verbatim description including the interpretation of the scale). Questions map to five moral foundations: Harm/Care, Fairness/Reciprocity, In-group/Loyalty, Authority/Respect, Purity/Sanctity. The results are typically presented as foundation-level scores, obtained by averaging the ratings of the questions associated with each foundation.
@@
Figure 3 reports foundation-level MFQ scores averaged over all models for different personas, and for the average persona. It gives an average characterization of the moral profile of models elicited by a given persona (see Appendix D for the persona descriptions). It is intereting to compare the average score across models with no persona-conditioning with the average score across models over all personas, this gives a sense of the average moral response of models to persona role-play (see Appendix B for details). Complementary, in Table 1, we present the persona ID that attained the highest MFQ score averaged across models for each foundation. The full per-persona, per-model, and per-question MFQ ratings are available in our GitHub repository [costa2025llmms].
Table 1: Persona that attains the highest MFQ score averaged across models for each foundation.
@@
Mean Score 4.79 4.84 4.72 4.54 4.42
@@ 2.2 Experimental Methodology
  • Personas: We evaluate $|\mathcal{P}|=100$ persona descriptions drawn from prior work [ge2025scalingsyntheticdatacreation]. Full persona descriptions and the corresponding ID–description mappings are provided in Appendix D.
  • Prompting: For each persona and question, the model receives a role-playing instruction: “You are roleplaying as the following persona:", followed by the persona description text and one of the $|\mathcal{Q}|=30$ MFQ questions. [^1] We query one MFQ question at a time rather than the full questionnaire in a single prompt to avoid sequence- and order-dependent effects. We instruct the models to start their response with the rating (an integer from 0 to 5), followed by their reasoning. Exact prompt templates are provided in Appendix A.
@@
  • Models: We included: Claude Haiku 4.5, Claude Sonnet 4.5, DeepSeek V3,DeepSeek V3.1, Gemini 2.5 Flash Lite, Gemini 2.5 Flash, GPT-4.1, GPT-4.1 Mini, GPT-4.1 Nano, GPT-4o, GPT-4o Mini, GPT-5, GPT-5 Mini, GPT-5 Nano, Grok 4 and Grok 4 Fast.
  • Families: We group the above models in the following families: Claude, DeepSeek, Gemini, GPT-4, GPT-5 and Grok.
  • Logging: For each model we did a total of $|\mathcal{Q}|\times|\mathcal{P}|\times n=30\times 100\times 10=30,000$ requests. The resulting tables are available in our GitHub repository [costa2025llmms].
@@ 2.3 Statistical Analysis
$\bar{u}=\frac{1}{|\mathcal{P}|\,|\mathcal{Q}|}\sum_{p\in\mathcal{P}}\sum_{q\in\mathcal{Q}}u_{pq},\qquad\sigma_{\overline{u}}^{2}=\frac{1}{|\mathcal{P}||\mathcal{Q}|(|\mathcal{P}|\,|\mathcal{Q}|-1)}\sum_{p\in\mathcal{P}}\sum_{q\in\mathcal{Q}}(u_{pq}-\bar{u})^{2}.$ (2)
@@
$s_{qg}^{2}=\frac{1}{|\mathcal{P}_{g}|-1}\sum_{p\in\mathcal{P}_{g}}(\bar{y}_{pq}-\bar{y}_{gq})^{2},\qquad\bar{y}_{gq}=\frac{1}{|\mathcal{P}_{g}|}\sum_{p\in\mathcal{P}_{g}}\bar{y}_{pq}.$ (5)
@@
$\widetilde{S}=\frac{1}{G}\sum_{g=1}^{G}\widetilde{S}_{g},\qquad\widetilde{S}_{g}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}s_{qg},\qquad\sigma_{\widetilde{S}}=\sqrt{\frac{1}{G(G-1)}\sum_{g=1}^{G}(\widetilde{S}_{g}-\widetilde{S})^{2}},$ (6)
@@
To propagate uncertainty in both robustness and susceptibility we adopt a first-order approximation: we treat the cross-model averages $\mathbb{E}[\bar{u}]$ and $\mathbb{E}[\widetilde{S}$] as fixed constants and only propagate the uncertainty from $\widetilde{R}$ and $\widetilde{S}$. Linearizing Eqs. (4) and (7) around these values yields the closed-form standard errors $\sigma_{R}$ and $\sigma_{S}$ reported alongside each index. This analytical approximation was compared with a bootstrap approach, and both gave similar values.
@@ 3 Results
Our results for the overall moral robustness, Eq. (4), and moral susceptibility, Eq. (7), by model are displayed in Figure 4. For robustness, we see that model family explains most of the variance, with model size having no systematic effect. The Claude family is by a significant margin the most robust, while Grok are the least. At the model level Claude Sonnet 4.5 stand out as the most robust and GPT-5 Nano as the least. In contrast, moral susceptibility exhibits a mild family effect but a clear within-family size effect, with larger vairants being more susceptible. At the model level, Gemini 2.5 Flash is the most suscepbtile and GPT-5 Nano the least. Overall, the Grok family sits as the primary outlier, pairing comparatively low robustness with high susceptibility.
Figure 4: Left: moral robustness, Eq. (4): higher values indicate greater MFQ rating stability. Right: moral susceptibility, Eq. (7): higher values indicate larger persona-driven shifts in MFQ scores.
@@ 3.2 Moral Susceptibility
Our results for foundation-level moral susceptibility Eq. 7 are displayed in Figure 6. One can see that models have a more balanced susceptibility moral profile if compared with robustness, with no model scoring significantly higher across foundations. Interestingly, DeepSeek V3.1 and the Llama models have a more similar susceptibility profile, with low Harm/Care and Fairness/Reciprocity, in comparison with the other foundations. In contrast, Gemini-2.5 Flahs Lite, GPT-4.1 Nano, GPT-5 Nano, have a high Harm/Care and Fairness/Reciprocity.
@@ 3.3 Correlation Between Robustness and Susceptibility
Table 3 summarises the Pearson correlations from Eq. (9) at the model and family levels. With all models included we see a positive association between robustness and susceptibility that is more pronounced at the family level, with Fairness/Reciprocity and Harm/Care showing the strongest dependencies while Purity/Sanctity remains negative. The Grok family suppresses these values because of its low robustness and high susceptibility, particularly within Purity/Sanctity. Excluding Grok lifts every foundation, reinforcing the positive trends without altering the qualitative ordering across foundations.
Table 3: Pearson correlation between robustness and susceptibility overall and by foundation. Columns on the right report the same metrics after excluding the Grok family.
All models Excluding Grok
Foundation Model $r_{RS}$ Family $\overline{r}_{RS}$ Model $r_{RS}$ Family $\overline{r}_{RS}$
All foundations $0.27\pm 0.07 +$$0.29\pm 0.09 +$$0.41\pm 0.07 +$$0.48\pm 0.09$
Authority/Respect $0.07\pm 0.09 +$$0.12\pm 0.14 +$$0.20\pm 0.10 +$$0.33\pm 0.16$
Fairness/Reciprocity $0.36\pm 0.08 +$$0.46\pm 0.10 +$$0.45\pm 0.08 +$$0.62\pm 0.09$
Harm/Care $0.26\pm 0.06 +$$0.39\pm 0.07 +$$0.38\pm 0.06 +$$0.59\pm 0.07$
In-group/Loyalty $0.25\pm 0.07 +$$0.41\pm 0.10 +$$0.31\pm 0.07 +$$0.50\pm 0.10$
Purity/Sanctity $0.18\pm 0.09$ $0.24\pm 0.11$ $0.10\pm 0.09$ $0.15\pm 0.12$
@@ Appendix A Prompt Format
Our experiment prompt includes: a persona role-play instruction, one MFQ relevance or agreement question, and a leading integer response instruction. The question texts are sourced from the Moral Foundations Questionnaire [moralfoundations2017questionnaires] and are not repeated here. Below we show the exact prompt formats used.
@@ Appendix B Moral Foundation Tables
1 $4.49\pm 0.05$ $4.61\pm 0.05$ $3.75\pm 0.14$ $2.87\pm 0.20$ $2.44\pm 0.23$
8 $3.73\pm 0.12$ $3.65\pm 0.10$ $3.58\pm 0.08$ $3.43\pm 0.07$ $2.31\pm 0.10$
16 $3.84\pm 0.14$ $3.92\pm 0.14$ $2.28\pm 0.13$ $2.09\pm 0.14$ $1.98\pm 0.17$
29 $3.78\pm 0.10$ $3.96\pm 0.09$ $2.59\pm 0.12$ $2.32\pm 0.12$ $1.59\pm 0.15$
30 $4.06\pm 0.09$ $4.68\pm 0.06$ $2.61\pm 0.14$ $2.80\pm 0.10$ $1.86\pm 0.15$
33 $4.26\pm 0.07$ $4.28\pm 0.07$ $3.05\pm 0.11$ $2.47\pm 0.16$ $2.33\pm 0.16$
47 $4.50\pm 0.07$ $52\pm 0.07 4.4$$44\pm 0.08$ $4.05\pm 0.11$ $3.87\pm 0.09$
60 $18\pm 0.11 3.3$$33\pm 0.10$ $3.65\pm 0.09$ $3.23\pm 0.08$ $1.88\pm 0.15$
69 $4.20\pm 0.05$ $4.27\pm 0.07$ $3.18\pm 0.11$ $2.61\pm 0.14$ $2.09\pm 0.18$
70 $4.47\pm 0.10$ $4.29\pm 0.11$ $2.58\pm 0.08$ $2.34\pm 0.08$ $2.40\pm 0.08$
74 $4.55\pm 0.07$ $4.60\pm 0.07$ $4.00\pm 0.06$ $60\pm 0.07 3.3$$35\pm 0.09$
75 $4.18\pm 0.11$ $4.68\pm 0.06$ $4.72\pm 0.07$ $3.23\pm 0.19$ $3.06\pm 0.14$
77 $42\pm 0.07 4.4$$49\pm 0.07$ $3.26\pm 0.12$ $2.92\pm 0.11$ $2.69\pm 0.14$
80 $4.41\pm 0.11$ $4.69\pm 0.08$ $2.49\pm 0.13$ $2.40\pm 0.10$ $1.61\pm 0.12$
Average $4.17\pm 0.01$ $4.25\pm 0.01$ $3.16\pm 0.02$ $2.83\pm 0.02$ $2.40\pm 0.02$
It is intereting to compare the One can see that the only foundation for which models performed better than the average persona was Farirness/Reciprocity with the average over models being $4.36\pm 0.07$ and the average over personas being $4.25\pm 0.01$. In contrast, the one for which the average over personas has the largest relative difference is Purity/Sanctity, $1.95\pm 0.11$ vs $2.40\pm 0.02$.

### v2

# arXiv:2511.08565v2[cs.CL] 27 Feb 2026

@@ Abstract
Large Language Models, Moral Psychology, Persona Role-Play
@@ 1 Introduction
As large language models (LLMs) move into interactive, multi-agent settings, reliable benchmarks for their social reasoning are essential. Recent evaluations probe theory-of-mind, multi-agent interactions under asymmetric information, cooperation, and deception through controlled role-play and game-theoretic tasks (Zhou and others, 2024; Pan and others, 2023; Bianchi and others, 2024; Chen and others, 2024; Costa and Vicente, 2025). Complementary datasets benchmark social commonsense, moral judgment, and self-recognition capabilities (Sap and others, 2019; Hendrycks and others, 2021; Bai et al., 2025b). Motivated by this landscape, we focus on moral judgment as a core facet of social decision-making and alignment.
This paper introduces a benchmark that combines persona role-play—prompting a LLM to assume a specific character—with the Moral Foundations Questionnaire (MFQ, 2017), a widely used instrument in moral psychology that measures five moral foundations: Harm/Care, Fairness/Reciprocity, In-group/Loyalty, Authority/Respect, and Purity/Sanctity (Graham et al., 2009; Haidt and Graham, 2007; MFQ, 2017). We elicit LLMs to respond to the MFQ while role-playing personas drawn from Ge et al. (2025). From these responses, we define two complementary quantities: moral robustness, the stability of MFQ scores over personas under repeated sampling, and moral susceptibility, the sensitivity of MFQ scores to persona variation. See Fig. 1 for a conceptual overview diagram. These metrics are defined in Eq. (4) and Eq. (8), each with foundation-level decompositions and uncertainty estimates.
@@
Applying this framework across contemporary model families and sizes, we find that model family accounts for most of the variance in moral robustness, with no systematic effect of model size. In contrast, moral susceptibility shows a mild family effect but a clear within-family size trend, with larger variants being more susceptible. Among individual models, Claude 4.5 Sonnet is the most robust and Grok 4 Fast the least. Conversely, Gemini 2.5 Flash is the most susceptible, while Llama-4 Scout is the least. Overall, we observe a a positive correlation between robustness and susceptibility, which is more pronounced at the family level, as seen in Section 3.3.
Recent research has examined the moral and social behavior of LLMs through the lens of the MFQ, exploring their value orientations, cultural variability, and alignment with human moral judgments (Abdulhai et al., 2024; Nunes et al., 2024; Aksoy, 2024; Bajpai et al., 2024; Ji et al., 2025). Parallel efforts study persona role-playing as a mechanism for conditioning model behavior, including benchmarks, interactive environments, and diagnostic analyses (Tseng et al., 2024; Wang et al., 2023; Samuel et al., 2025; Yu et al., 2025; Xu et al., 2024; Boudouri et al., 2025; Bai et al., 2025a). Our MFQ persona framework bridges these directions by systematically quantifying how persona conditioning alters moral judgments, separating the effects of repeated sampling (moral robustness) from those of persona variation (moral susceptibility). In addition, we report MFQ profiles for both unconditioned and persona-conditioned settings, providing a comparative view of baseline moral tendencies and persona-driven moral shifts across models.
@@ 2.1 Moral Foundation Questionnaire
The Moral Foundations Questionnaire (MFQ, 2017) is a widely used instrument in moral psychology (Graham et al., 2009; Haidt and Graham, 2007; MFQ, 2017) and comprises 30 questions split into two sections. The first includes 15 relevance judgments, which assess how relevant certain considerations are when deciding what is right or wrong, and the second includes 15 agreement statements, which measure the level of agreement with specific moral propositions (Graham et al., 2011; MFQ, 2017). In both sections, respondents answer each item using an integer scale from 0 to 5, representing in the first section the perceived relevance of the consideration and in the second the degree of agreement with the statement (see Appendix A for a verbatim description including the interpretation of the scale). Questions map to five moral foundations: Harm/Care, Fairness/Reciprocity, In-group/Loyalty, Authority/Respect, Purity/Sanctity. The results are typically presented as foundation-level scores, obtained by averaging the ratings of the questions associated with each foundation.
@@
Figure 3 reports foundation-level MFQ scores averaged over all models for different personas, and for the average persona. It gives an average characterization of the moral profile of models elicited by a given persona (see Appendix D for the persona descriptions). It is intereting to compare the average score across models with no persona-conditioning with the average score across models over all personas, this gives a sense of the average moral response of models to persona role-play (see Appendix B for details). Complementary, in Table 1, we present the persona IDs that attained the highest and lowest MFQ scores averaged across models for each foundation. The full per-persona, per-model, and per-question MFQ ratings will be made available online if the paper is accepted.
Table 1: Personas that attain the highest and lowest MFQ scores averaged across models for each foundation. See Appendix D for the persona descriptions.
@@
Highest Mean Score 4.79 4.84 4.72 4.54 4.42
Persona ID 86 58 86 13 86
Lowest Mean Score 2.94 3.24 1.73 1.64 1.15
@@ 2.2 Experimental Methodology
  • Personas: We evaluate $|\mathcal{P}|=100$ persona descriptions drawn from prior work (Ge et al., 2025). Full persona descriptions and the corresponding ID–description mappings are provided in Appendix D.
  • Prompting: For each persona and question, the model receives a role-playing instruction: “You are roleplaying as the following persona:", followed by the persona description text and one of the $|\mathcal{Q}|=30$ MFQ questions. [^1] We query one MFQ question at a time rather than the full questionnaire in a single prompt to avoid sequence- and order-dependent effects. We instruct the models to start their response with the rating (an integer from 0 to 5), followed by their reasoning. Exact prompt templates are provided in Appendix A. We use decoding temperature $T=0.1$ to keep stochasticity consistent across models.
@@
  • Models: We included: Claude Haiku 4.5, Claude Sonnet 4.5, DeepSeek V3,DeepSeek V3.1, Gemini 2.5 Flash Lite, Gemini 2.5 Flash, GPT-4.1, GPT-4.1 Mini, GPT-4.1 Nano, GPT-4o, GPT-4o Mini, Grok 4 and Grok 4 Fast.
  • Families: We group the above models in the following families: Claude, DeepSeek, Gemini, GPT-4, and Grok.
  • Logging: For each model we did a total of $|\mathcal{Q}|\times|\mathcal{P}|\times n=30\times 100\times 10=30,000$ requests. The resulting tables will be made available online if the paper is accepted.
@@ 2.3 Statistical Analysis
$\bar{u}$ $=\frac{1}{|\mathcal{P}|\,|\mathcal{Q}|}\sum_{p\in\mathcal{P}}\sum_{q\in\mathcal{Q}}u_{pq},$ (2)
$\sigma_{\overline{u}}^{2}$ $=\frac{1}{|\mathcal{P}||\mathcal{Q}|(|\mathcal{P}|\,|\mathcal{Q}|-1)}\sum_{p\in\mathcal{P}}\sum_{q\in\mathcal{Q}}(u_{pq}-\bar{u})^{2}.$
@@
$s_{qg}^{2}$ $=\frac{1}{|\mathcal{P}_{g}|-1}\sum_{p\in\mathcal{P}_{g}}(\bar{y}_{pq}-\bar{y}_{gq})^{2},$ (5)
$\bar{y}_{gq}$ $=\frac{1}{|\mathcal{P}_{g}|}\sum_{p\in\mathcal{P}_{g}}\bar{y}_{pq}.$ (6)
@@
$\widetilde{S}$ $=\frac{1}{G}\sum_{g=1}^{G}\Big(\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}s_{qg}\Big),$ (7)
$\sigma_{\widetilde{S}}$ $=\sqrt{\frac{1}{G(G-1)}\sum_{g=1}^{G}(\widetilde{S}_{g}-\widetilde{S})^{2}}.$
@@
To propagate uncertainty in both robustness and susceptibility we adopt a first-order approximation: we treat the cross-model averages $\mathbb{E}[\bar{R}]$ and $\mathbb{E}[\widetilde{S}$] as fixed constants and only propagate the uncertainty from $\widetilde{R}$ and $\widetilde{S}$. Linearizing Eqs. (4) and (8) around these values yields the closed-form standard errors $\sigma_{R}$ and $\sigma_{S}$ reported alongside each index. This analytical approximation was compared with a bootstrap approach, and both gave similar values.
@@
2.4 Temperature
We fix the decoding temperature to $T=0.1$ for every model. Temperature directly modulates the dispersion terms we estimate: in the deterministic $T=0$ limit all repetitions coincide, yielding $\bar{u}=0$ and driving $R$ to infinity, while larger temperatures inflate both within- and across-persona variances $u_{pq}$ and $s_{qg}$. Consequently, comparing models requires that every run uses the same temperature.
@@ 3 Results
Our results for the overall moral robustness, Eq. (4), and moral susceptibility, Eq. (8), by model are displayed in Figures 4 and 5. Figure 4 shows that model family explains most of the robustness variance, with model size having no systematic effect. The Claude family is by a significant margin the most robust, while Grok are the least. At the model level Claude Sonnet 4.5 stands out as the most robust and Grok-4 as the least. In contrast, Figure 5 highlights that moral susceptibility exhibits a mild family effect but a clear within-family size effect, with larger variants being more susceptible. At the model level, Gemini 2.5 Flash is the most susceptible and Llama-4 Scout as the least. Overall, the Grok family sits as the primary outlier, pairing comparatively low robustness with high susceptibility across both panels.
Figure 4: Moral robustness, Eq. (4): higher values indicate greater MFQ rating stability.
Figure 5: Moral susceptibility, Eq. (8): higher values indicate larger persona-driven shifts in MFQ scores.
@@ 3.2 Moral Susceptibility
Our results for foundation-level moral susceptibility Eq. 8 are displayed in Figure 7. One can see that models have a more balanced susceptibility moral profile if compared with robustness, with no model scoring significantly higher across foundations. Interestingly, DeepSeek V3.1 and the Llama models have a more similar susceptibility profile, with low Harm/Care and Fairness/Reciprocity, in comparison with the other foundations. In contrast, Gemini-2.5 Flahs Lite, GPT-4.1 Nano, have a high Harm/Care and Fairness/Reciprocity.
@@ 3.3 Correlation Between Robustness and Susceptibility
Table 3 summarises the Pearson correlations from Eq. (10) at the model and family levels. With all models included we see a positive association between robustness and susceptibility that is more pronounced at the family level, with Fairness/Reciprocity and Harm/Care showing the strongest dependencies while Purity/Sanctity remains negative. The Grok family suppresses these values because of its low robustness and high susceptibility, particularly within Purity/Sanctity; removing Grok (figures omitted for space) raises every coefficient but leaves the qualitative ordering unchanged.
Table 3: Pearson correlation between robustness and susceptibility overall and by foundation.
Foundation Model $r_{RS}$ Family $\overline{r}_{RS}$
All foundations $0.27\pm 0.07 +$$0.29\pm 0.09$
Authority/Respect $0.07\pm 0.09 +$$0.12\pm 0.14$
Fairness/Reciprocity $0.36\pm 0.08 +$$0.46\pm 0.10$
Harm/Care $0.26\pm 0.06 +$$0.39\pm 0.07$
In-group/Loyalty $0.25\pm 0.07 +$$0.41\pm 0.10$
Purity/Sanctity $0.18\pm 0.09$ $0.24\pm 0.11$
@@ Acknowledgments
Impact Statements
This paper presents work whose goal is to advance the understanding of moral behavior in large language models by systematically analyzing how moral judgments change under persona role-play. By introducing a benchmark grounded in moral psychology, this work may contribute to improved evaluation and design of more reliable and value-aligned AI systems. At the same time, insights into moral susceptibility could potentially be misused to deliberately influence or manipulate model outputs in sensitive contexts. We view this research as primarily foundational and methodological, and we do not anticipate significant negative societal consequences beyond those already well established for research on model alignment and controllability.
@@ Appendix A Prompt Format
Our experiment prompt includes: a persona role-play instruction, one MFQ relevance or agreement question, and a leading integer response instruction. The question texts are sourced from the Moral Foundations Questionnaire (MFQ, 2017) and are not repeated here. Below we show the exact prompt formats used.
@@ Appendix B Moral Foundation Tables
1 $4.46\pm 0.06$ $4.60\pm 0.07$ $3.76\pm 0.16$ $2.80\pm 0.21$ $2.31\pm 0.26$
8 $3.66\pm 0.15$ $3.55\pm 0.13$ $57\pm 0.10 3.3$$38\pm 0.08$ $2.21\pm 0.11$
16 $3.75\pm 0.18$ $3.81\pm 0.18$ $2.11\pm 0.12$ $1.92\pm 0.12$ $1.78\pm 0.18$
29 $3.69\pm 0.12$ $3.91\pm 0.11$ $54\pm 0.14 2.2$$25\pm 0.14$ $1.46\pm 0.15$
30 $4.09\pm 0.10$ $4.73\pm 0.05$ $2.46\pm 0.17$ $2.74\pm 0.13$ $1.81\pm 0.18$
33 $4.20\pm 0.08$ $4.23\pm 0.08$ $3.05\pm 0.13$ $34\pm 0.17 2.2$$27\pm 0.19$
47 $52\pm 0.09 4.4$$7\pm 0.10 4.4$$46\pm 0.10$ $4.03\pm 0.13$ $3.87\pm 0.11$
60 $3.05\pm 0.13$ $3.23\pm 0.12$ $3.68\pm 0.11$ $3.25\pm 0.10$ $1.81\pm 0.15$
69 $4.16\pm 0.06$ $4.25\pm 0.09$ $3.20\pm 0.14$ $2.57\pm 0.15$ $2.07\pm 0.23$
70 $4.56\pm 0.09$ $4.31\pm 0.11$ $2.55\pm 0.10$ $2.30\pm 0.11$ $2.39\pm 0.11$
74 $4.57\pm 0.08$ $4.60\pm 0.08$ $4.02\pm 0.08$ $67\pm 0.09 3.3$$35\pm 0.13$
75 $4.13\pm 0.14$ $4.64\pm 0.08$ $4.83\pm 0.05$ $3.26\pm 0.23$ $3.09\pm 0.17$
77 $41\pm 0.08 4.4$$47\pm 0.08$ $3.18\pm 0.14$ $2.83\pm 0.11$ $2.60\pm 0.17$
80 $4.44\pm 0.11$ $4.71\pm 0.08$ $2.37\pm 0.14$ $2.30\pm 0.11$ $1.51\pm 0.15$
It is interesting to compare the average score across models with no persona-conditioning to the average over all personas; the only foundation where models exceed the persona average is Fairness/Reciprocity with the average over models being $4.36\pm 0.07$ and the average over personas being $4.25\pm 0.01$. In contrast, the one for which the average over personas has the largest relative difference is Purity/Sanctity, $1.95\pm 0.11$ vs $2.40\pm 0.02$.